OpenStudy (anonymous):

evaluate the lim lim x--> pi/4 sinx/x

OpenStudy (nenadmatematika):

=sin(pi/4)/pi/4=(1/sqrt2)/(pi/4)=4/(pi*sqrt2)

OpenStudy (anonymous):

the andswer i got was 2sqrt2/pi

OpenStudy (anonymous):

answer***

OpenStudy (nenadmatematika):

which is the same :D

OpenStudy (anonymous):

ok thanks